OpenStudy (anonymous):

1. An astronaut takes an object to the moon where there is less gravity. how does the mass and weight of the object on the moon would compare to its mass and weight on Earth?

OpenStudy (sidsiddhartha):

mass of a object is a fixed property ,an objects mass stays the same every where so The objects mass will never change. but weight of an object is force which is equal to \[F=mg\] Earth's acceleration of gravity is 9.81 m/s^2, and the Moons acceleration of gravity is 1.63 m/s^2 - which is exactly 1/6 of Earth's gravity. so take an example-- if an object had a mass of 20 kilograms, its weight would be 196.2 Newtons \[W_{earth}=20*9.81=196.2\] but for moon it will be \[W_{moon}=20*\frac{ 9.81 }{ 6 }=32.7\] hope it helps @cosplay_queen